40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  TCS elog  Not logged in ELOG logo
Message ID: 151     Entry time: Mon May 30 10:53:00 2011
Author: Aidan 
Type: Laser 
Category: Hartmann sensor 
Subject: Defocus vs time 

 I've had the output from a fiber projected about 400mm onto the Hartmann sensor for around 5 days now. (The divergence angle from the fiber is around 86 mrad).

I played around with the temperature of the lab to induce some defocus changes in the Hartmann sensor. The system is mostly linear, but there are relatively frequent jumps in the defocus of the order of 1E-4 m^-1. This may be due to a number of things - the Hartmann plate may be moving, the fiber holder may be shifting back and forth, there may be some issue with the source wavelength shifting.

  • A change in defocus, dS, of around 1E-4 m^-1 from a point source at 400mm from the Hartmann plate (dS/dx = -1/x^2), corresponds to a change in the end position of the fiber of around 16 microns. Seems a little big ... unless it's not secured properly ...
  • Also the defocus vs temperature slope is different from what James measured last year. There is an additional -4E-5 m^-1 K^-1 due to the expansion of the stainless steel table moving the point source further or closer to the Hartmann plate. That leaves about -3E-4 m^-1 K^-1 for the Hartmann sensor. This is roughly twice what James measured last year. 

Sun 30th May 2011 - 11:40AM - the z-axis control on the NewFocus 9091 fiber coupling mount was not tightened. I tightened that to secure the control.

Attachment 1: HWS_defocus_vs_temperature.pdf  46 kB  Uploaded Mon May 30 11:53:23 2011  | Hide | Hide all
HWS_defocus_vs_temperature.pdf
ELOG V3.1.3-